I am trying to simplify the following:
(x^2)^5 * x / x^3
I got to this:
x^2.5 / x^3 * x / x^3 = x^10 / x^3 * x / x^3
Now I am stumped.

Answer by edjones(8007)

(x^2)^5 * x / x^3
Let's do it in steps.
(x^2)^5=x^10 multiply when exp^exp.
x^10*x=x^11 add exp when multiplying.
x^11/x^3=x^8 subtract exp when dividing. Ans.
It's easy when you know how.
Ed